# Implicit Bias in Matrix Factorization and its Explicit Realization in a New Architecture

### v1

@@ Abstract
Gradient descent for matrix factorization is known to exhibit an implicit bias toward approximately low-rank solutions. While existing theories often assume the boundedness of iterates, empirically the bias persists even with unbounded sequences. We thus hypothesize that implicit bias is driven by divergent dynamics markedly different from the convergent dynamics for data fitting. Using this perspective, we introduce a new factorization model: $X\approx UDV^{\top}$, where $U$ and $V$ are constrained within norm balls, while $D$ is a diagonal factor allowing the model to span the entire search space. Our experiments reveal that this model exhibits a strong implicit bias regardless of initialization and step size, yielding truly (rather than approximately) low-rank solutions. Furthermore, drawing parallels between matrix factorization and neural networks, we propose a novel neural network model featuring constrained layers and diagonal components. This model achieves strong performance across various regression and classification tasks while finding low-rank solutions, resulting in efficient and lightweight networks.
@@ 1 Introduction
The Burer–Monteiro (BM) factorization (Burer & Monteiro, 2003) is a classical technique for obtaining low-rank solutions in optimization. It can be viewed as a simple neural network that uses a single layer of hidden neurons under linear activation. Indeed, say one has the factorization $X=UV^{T}$ where $U\in\mathbb{R}^{d\times r}$ and $V\in\mathbb{R}^{c\times r}$, then $U$ and $V$ can be thought of as the weights of the first and second layers, while $r$ represents the number of hidden neurons. However, despite the similarity suggested by this view, there is a clear distinction between BM factorization and neural networks about how the rank $r$ is chosen. In BM factorization, $r$ is typically chosen to be small, close to the rank of the desired solution. Neural networks, on the other hand, often succeed even in overparametrized settings where $r$ is large.
Recent findings of implicit regularization in matrix factorization narrows the gap between these two perspectives. For instance, in an effort to explain the empirical success of overparametrized neural networks, Gunasekar et al. (2017) demonstrate that gradient descent (with certain parameter selection) on BM factorization tends to converge toward approximately low-rank solutions even when $r=d$. Based on this observation, they conjecture that “with small enough step sizes and initialization close enough to the origin, gradient descent on full-dimensional factorization converges to the minimum nuclear norm solution.”
In a follow-up work, however, Razin & Cohen (2020) presented a counter-example demonstrating that implicit regularization in BM factorization cannot be explained by minimal nuclear norm, or in fact any norm. Specifically, they showed that there are instances where the gradient method applied to BM factorization yields a diverging sequence, and all norms thus grow toward infinity. Intriguingly, despite this divergence, they found that the rank of the estimate decreases toward its minimum.
Although this phenomenon might seem surprising initially, it is not uncommon for diverging sequences to follow a structured path. A prime example is the Power Method, the fundamental algorithm for finding the largest eigenvalue and eigenvector pair of a matrix. Starting from a random initial point $x_{0}$, the Power Method iteratively updates the estimate by multiplying it with the matrix. This process amplifies the component of the vector that aligns with the direction of the dominant eigenvector more than the other components, progressively leading $x_{k}$ to align with this eigenvector. In practical implementations, $x_{k}$ is scaled after each iteration to avoid numerical issues from divergence.
This perspective serves as the foundational motivation guiding our approach. Specifically, our key insight is that the implicit regularization in BM factorization (and neural networks) is driven by divergent dynamical behavior. This is markedly different from the standard (convergent) optimization dynamics helping with the data fitting. In this context, we hypothesize that these forces do not merely coexist but actively compete, influencing model behavior and performance in fundamentally conflicting ways. Our main goal in the development of this paper is to devise an approach that unravels these competing forces.
@@ 1.1 Overview of main contributions
    A novel formulation for matrix factorization. We model $X=UDV^{\top}$, where $U$ and $V$ are constrained within Frobenius norm balls. Projection onto this ball results in a scaling step similar to the Power Method. The middle term $D$ is a diagonal matrix that allows the model to explore the entire search space despite $U$ and $V$ being bounded.
    Empirically we demonstrate that the gradient method applied to the proposed formulation exhibits a pronounced implicit bias toward low-rank solutions. We compare our formulation against standard BM factorization with two unconstrained factors. Specifically, we investigate key factors such as step size and initialization, which prior work suggests might be contributing to implicit bias. We find that our factorization approach largely obviates the need to rely on these conditions: it consistently finds truly (rather than approximately) low-rank solutions across a wide range of initializations and step-sizes in our experiments. We believe these findings should be of broader interest to research on implicit bias.
@@
    A novel neural network architecture. Motivated by the strong bias for low-rank solutions of the proposed factorization, we subsequently extend it to deep neural networks. Specifically, we do so by adding constrained layers and diagonal components. We demonstrate numerically that this constrained model performs on par with, or even better than, the standard architecture across various regression and classification tasks. Importantly, even here, our approach exhibits bias towards low-rank solutions, resulting in a natural pruning procedure to obtain compact, lightweight networks without compromising performance.
2 Related Work
Burer-Monteiro factorization. BM factorization is initially proposed for solving semi-definite programs (Burer & Monteiro, 2003, 2005; Boumal et al., 2016, 2020; Cifuentes, 2021) and is recognized for its efficiency in addressing low-rank matrix optimization problems, see (Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017, 2018; Hsieh et al., 2018; Sahin et al., 2019; Lee et al., 2022; Yalçın et al., 2023) and the references therein. Building on the connections between training problems for (non-linear) two-layer neural networks and convex (copositive) formulations (see (Pilanci & Ergen, 2020; Ergen & Pilanci, 2020; Sahiner et al., 2021) and the references therein), Sahiner et al. (2023) recently introduced BM factorization to solve convex formulations for various neural network architectures, including fully connected networks with ReLU and gated ReLU (Fiat et al., 2019) activations, two-layer convolutional neural networks (LeCun et al., 1989a), and self-attention mechanisms based on transformers (Vaswani et al., 2017).
@@ 2 Related Work
Implicit regularization. While neural network-based systems are rapidly emerging as a dominant technology, the mechanisms underlying their generalization capabilities are yet to be fully understood. One promising line of research aims to explain this success through the concept of ‘implicit regularization,’ which is induced by the optimization methods and formulations used during neural network training (Neyshabur et al., 2014, 2017; Neyshabur, 2017). Several studies have explored matrix factorization to investigate implicit bias in linear neural networks (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021). Le & Jegelka (2022) extended these results to the final linear layers of nonlinear ReLU-activated feedforward networks with fully connected layers and skip connections. More recently Timor et al. (2023) investigated implicit regularization in ReLU networks. Much of the existing work focuses on gradient flow dynamics in the limit of infinitesimal learning rates. In particular, Gidel et al. (2019) examined discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
@@
Constrained neural networks. Regularizers are frequently used in neural network training to prevent overfitting and improve generalization (Goodfellow et al., 2016), or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017). However, it is conventional to apply these regularizers as penalty functions in the objective rather than as hard constraints, addressing them within gradient-based optimization via their (sub)gradients. This approach is likely favored due to the ease of implementation, as pre-built functions are readily available in common neural network packages. Several independent studies have also applied proximal methods across different networks and applications, finding that proximal-based methods tend to yield solutions with more pronounced structures (Bai et al., 2019; Yang et al., 2020; Yun et al., 2021; Yurtsever et al., 2021; Yang et al., 2022). Structural regularization in the form of hard constraints, however, appears to be rare in neural network training. One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022), where constraints are necessary due to the algorithm’s requirement for a bounded domain.
There are other use cases of constraints in neural networks. For instance, constraints can be applied to ensure adherence to real-world conditions in applications where they are necessary (Pathak et al., 2015; Márquez-Neila et al., 2017; Jia et al., 2017; Kervadec et al., 2019; Weber et al., 2021), or to incorporate physical laws in physics-informed neural networks (Raissi et al., 2019; Lu et al., 2021; Patel et al., 2022). In these cases, the feasible set is typically complex and difficult to project onto; therefore, optimization algorithms like primal-dual methods or augmented Lagrangian techniques are used. Constraints are also present in lifted neural networks, a framework where the training problem is reformulated in a higher-dimensional ‘lifted’ space. In this space, conventional activation functions like ReLU, used in the original formulation, are expressed as constraints (Askari et al., 2018; Sahiner et al., 2021; Bartan & Pilanci, 2021; Prakhya et al., 2024).
Pruning. Neural networks are often trained in an over-parameterized regime to enhance generalization and avoid getting stuck in poor local minima. However, these models can suffer from excessive memory and computational resource demands, making them less efficient for deployment in real-world applications (Chang et al., 2021). Several compression techniques have been studied in the literature, including parameter quantization (Krishnamoorthi, 2018), knowledge distillation (Gou et al., 2021), and pruning.
Pruning reduces the number of parameters, resulting in more compact and efficient models that are easier to deploy. The literature on pruning is extensive, with diverse methods proposed with distinct characteristics. Various criteria are used to determine which weights to prune, including second-order derivative-based methods (LeCun et al., 1989b; Hassibi & Stork, 1992), magnitude-based pruning (Janowsky, 1989; Han et al., 2015), saliency heuristics (Mozer & Smolensky, 1988; Lee et al., 2018), and matrix or tensor factorization-based techniques (Xue et al., 2013; Sainath et al., 2013; Jaderberg et al., 2014; Lebedev et al., 2015; Swaminathan et al., 2020), among others. A comprehensive review of pruning methods is beyond the scope of this paper due to space limitations and the diversity of approaches. For more detailed reviews, we refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024), and the references therein.
@@ 3 Matrix Factorization with a Diagonal Component
Consider the matrix sensing problem, where the goal is to recover a positive semidefinite (PSD) matrix $X\in\mathbb{S}_{+}^{d\times d}$ from a set of linear measurements $\smash{b=\mathcal{A}(X)\in\mathbb{R}^{n}}$. We define $\smash{\mathcal{A}:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}^{n}}$ through symmetric measurement matrices $\smash{A_{1},\ldots,A_{n}\in\mathbb{S}^{d\times d}}$, such that $\mathcal{A}(X)=[\langle A_{1},X\rangle~{}\cdots~{}\langle A_{n},X\rangle]^{\top}$ and $\mathcal{A}^{\top}y=\sum_{i=1}^{n}y_{i}A_{i}.$ We particularly focus on the data-scarce setting where $n\ll d^{2}$. A popular example is the matrix completion problem, where the goal is to reconstruct a matrix $X$ from a subset of its entries. This problem is inherently under-determined; however, successful recovery is possible if $X$ is known to be low-rank (Candes & Recht, 2012).
Remark 1.
We focus on the recovery problem of a PSD matrix for simplicity; note that a non-square matrix sensing problem can also be reformulated as a PSD matrix sensing problem through a simple transformation (Park et al., 2017).
The problem described above can be cast as the following rank-constrained optimization problem:
$\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad f(X):=\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{subj.\ to}\quad\mathrm{rank}(X)\leq r.$ (1)
Although rank-constrained matrix optimization problems are typically NP-hard, various methods have been developed to provide practical approximations. Examples include hard thresholding algorithms (Jain et al., 2010; Goldfarb & Ma, 2011; Kyrillidis & Cevher, 2014), convex relaxation techniques (Candes & Recht, 2012; Recht et al., 2010), and BM factorization (Burer & Monteiro, 2003; Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017).
The main idea behind BM factorization is to reparametrize the decision variable $X$ as $UU^{\top}$, where the factor $U\in\mathbb{R}^{d\times r}$, and $r$ is a positive integer that controls the rank of the resulting product. Problem (1) can then be reformulated as:
@@
Despite the fact that finding the global minimum of (2) remains challenging, a local solution can be approximated using the gradient method (Lee et al., 2016). Initializing at $U_{0}\in\mathbb{R}^{d\times r}$, perform the iteration:
$U_{k+1}=U_{k}-\eta\nabla_{U}f(U_{k}U_{k}^{\top}),$ (3)
where $\eta>0$ is the step-size, and the gradient $\nabla_{U}f$ is computed as
$\nabla_{U}f(UU^{\top})=2\nabla f(UU^{\top})U=2\mathcal{A}^{\top}(\mathcal{A}(UU^{\top})-b)U.$
Selecting the factorization rank $r$ is a critical decision that might impact the solution. A small $r$ may lead to spurious local minima, resulting in inaccurate outcomes (Waldspurger & Waters, 2020). Conversely, a large $r$ might weaken rank regularization, rendering the problem underdetermined. Conventional wisdom in BM factorization suggests finding a moderate choice that balances these two extremes. However, a key observation in (Gunasekar et al., 2017) is that the gradient method applied to (2) exhibits a tendency towards approximately low-rank solutions (i.e., a decaying singular value spectrum) even when $r=d$. Below, we restate their conjecture:
Conjecture in (Gunasekar et al., 2017). Suppose gradient flow (i.e., gradient descent with an infinitesimally small step-size) is initialized at a full-rank matrix arbitrarily close to the origin. If the limit of the gradient flow, $X_{\text{GF}}=UU^{\top}$, exists and is a global optimum of (1) with $\mathcal{A}(X_{\text{GF}})=b$, then $X_{\text{GF}}$ is the minimal nuclear-norm solution to (1).
Remark 2.
This conjecture outlines three conditions for implicit bias: a small step-size, near-origin full-rank initialization, and consistent measurements ($\mathcal{A},b$). However, these conditions are often unconventional. For example, small step-sizes are typically avoided because they slow down convergence. Similarly, near-origin initialization is counterintuitive, as the origin is a trivial spurious stationary point of (2).
@@ 3.1 The Proposed Factorization
$\min_{\begin{subarray}{c}U\in\mathbb{R}^{d\times r}\\ D\in\mathbb{R}^{r\times r}\end{subarray}}~{}\frac{1}{2}\|\mathcal{A}(UDU^{\top})-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\|U\|_{F}\leq\alpha,~{}~{}D_{ii}\geq 0,~{}~{}D_{ij}=0,~{}~{}\forall i~{}\text{and}~{}\forall j\neq i,$ (4)
@@
Placing in multiple factors and with constraints, we perform projected-gradient updates on $U$ and $D$ with step-size $\eta>0$:
@@
3.2 Numerical Experiments on Matrix Factorization
We present numerical experiments comparing the empirical performance of the proposed approach with the classical BM factorization. Specifically, we examine the impact of initialization and step-size on the singular value spectrum of the resulting solution.
We set up a synthetic matrix completion problem to recover a PSD matrix $\smash{X_{\natural}=U_{\natural}U_{\natural}^{\top}}$, where the entries of $\smash{U_{\natural}\in\mathbb{R}^{100\times 3}}$ are independently and identically (iid) drawn from the standard Gaussian distribution. We randomly sample $n=900$ entries of $X_{\natural}$ and store them in the vector $b\in\mathbb{R}^{n}$. The goal is to recover $X_{\natural}$ from $b$ by solving problems (2) and (4). For initialization, we generate $U_{0}\in\mathbb{R}^{d\times d}$ with entries drawn iid from standard Gaussian distribution and rescale it to have the Frobenius norm $\xi>0$ (we investigate the impact of $\xi$). We initiate $D_{0}$ from the identity matrix.
The results are shown in Figure 1. First, we examine the impact of step-size. To this end, we fix $\xi=10^{-2}$ and test different values of $\eta$. In the left panel, we plot the objective residual as a function of iterations. As expected, we observe that a smaller step-size slows down convergence. In the right panel, we plot the singular value spectrum of the results attained after $10^{6}$ iterations. We observe no direct connection between step-size and implicit bias in BM factorization.
Impact of step-size ($\eta$), in noiseless setting, with fixed initialization.
Impact of initial distance to origin $(\xi)$, in noiseless setting, with fixed step-size.

### v2

@@ Abstract
Gradient descent for matrix factorization exhibits an implicit bias toward approximately low-rank solutions. While existing theories often assume the boundedness of iterates, empirically the bias persists even with unbounded sequences. This reflects a dynamic where factors develop low-rank structure while their magnitudes increase, tending to align with certain directions. To capture this behavior in a stable way, we introduce a new factorization model: $X\approx UDV^{\top}$, where $U$ and $V$ are constrained within norm balls, while $D$ is a diagonal factor allowing the model to span the entire search space. Experiments show that this model consistently exhibits a strong implicit bias, yielding truly (rather than approximately) low-rank solutions. Extending the idea to neural networks, we introduce a new model featuring constrained layers and diagonal components that achieves competitive performance on various regression and classification tasks while producing lightweight, low-rank representations.
@@ 1 Introduction
The Burer–Monteiro (BM) factorization (Burer & Monteiro, 2003) is a classical technique for obtaining low-rank solutions in optimization. One can view it as a simple neural network that uses a single layer of hidden neurons under linear activation. Given the factorization $X=UV^{\top}$ where $U\in\mathbb{R}^{d\times r}$ and $V\in\mathbb{R}^{c\times r}$, $U$ and $V$ can be interpreted as the weights of the first and second layers, and $r$ as the number of hidden neurons. As a result, matrix factorization has become a convenient model for investigating the principles underlying the empirical success of neural networks.
For instance, Gunasekar et al. (2017) demonstrate that gradient descent (with certain parameter selection) on BM factorization tends to converge toward approximately low-rank solutions even when $r=d$. Based on this observation, they conjecture that “with small enough step-sizes and initialization close enough to the origin, gradient descent on full-dimensional factorization converges to the minimum nuclear norm solution.”
In a follow-up work, Razin & Cohen (2020) present a counterexample demonstrating that implicit regularization in BM factorization cannot be explained by minimal nuclear norm, or in fact any norm. Specifically, they present instances where the gradient method applied to BM factorization yields a diverging sequence, and all norms thus grow toward infinity. Intriguingly, despite this divergence, they found that the rank of the estimate decreases toward its minimum.
In fact, it is not uncommon for diverging sequences to follow a structured path. A prime example is the Power Method, the fundamental algorithm for finding the largest eigenvalue and eigenvector pair of a matrix. Starting from a random initial point $x_{0}$, the Power Method iteratively updates the estimate by multiplying it with the matrix. This process amplifies the component of the vector that aligns with the direction of the dominant eigenvector more than the other components, progressively leading $x_{k}$ to align with this eigenvector. In practice, $x_{k}$ is scaled after each iteration to avoid numerical issues from divergence.
This perspective underpins our approach. We interpret implicit bias in BM factorization as arising from growth dynamics that amplify certain directions while suppressing others, thereby inducing low-rank structure. To capture this effect in a stable way, we propose a new factorization model that disentangles structured bias from instability.
@@ 1.1 Overview of main contributions
    A novel matrix factorization formulation. We introduce a new factorization model, $X=UDV^{\top}$, where $U$ and $V$ are constrained within Frobenius norm balls. Projection onto this ball results in a scaling step. The middle term $D$ is a diagonal matrix that allows the model to explore the entire search space despite $U$ and $V$ being bounded. [^1] The reader may notice a “syntactic” similarity with singular value decomposition; except using simpler Frobenius norm constraints on $U$ and $V$ instead of orthogonality.
    Through extensive empirics we demonstrate that the gradient method applied to the proposed formulation exhibits a pronounced implicit bias toward low-rank solutions. We compare our formulation against the standard BM factorization with two unconstrained factors. We investigate how step size and initialization influence implicit bias, building on prior work that points to these as potential contributors. We find that our factorization approach largely obviates the need to rely on these conditions: it consistently finds truly (rather than approximately) low-rank solutions across a wide range of initializations and step-sizes in our experiments.
@@
    A novel neural network architecture. Motivated by the strong bias for low-rank solutions of the proposed factorization, we extend it to deep neural networks. We do so by adding constrained layers and diagonal components. We show that this new model achieves competitive performance with standard architectures across various regression and classification tasks. Importantly, our approach exhibits bias towards low-rank solutions, resulting in a natural pruning procedure that delivers compact, lightweight networks without compromising performance.
1.2 Related Work
Burer-Monteiro factorization. BM factorization was proposed for solving semidefinite programs (Burer & Monteiro, 2003, 2005) and has been recognized for its efficiency in addressing low-rank optimization problems (Boumal et al., 2016; Park et al., 2018). Building on the connections between matrix factorization and training problems for two-layer neural networks, BM has served as foundational model for understanding implicit bias and developing theoretical insights.
@@ 1.2 Related Work
Implicit regularization. One promising line of research that aims to explain the successful generalization abilities of neural networks is that of ‘implicit regularization’ induced by the optimization methods and architectures (Neyshabur et al., 2014, 2017; Neyshabur, 2017). Several studies explore matrix factorization to investigate implicit bias (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021). Much of the existing work focuses on gradient flow dynamics in the limit of infinitesimal learning rates. Exceptionally, Gidel et al. (2019) examine discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
@@
Constrained neural networks. Regularizers are frequently used in neural network training to prevent overfitting and improve generalization, or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017). However, it is conventional to apply these regularizers as penalty functions in the objective rather than constraints. This approach is likely favored due to the ease of implementation, as pre-built functions are readily available in common neural network packages. Regularization in the form of constraints appears to be rare in neural network training. One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022). Recently, Pethick et al. (2025) revealed parallels between Frank-Wolfe on constrained networks and algorithms that post-process update steps, such as Muon (Jordan et al., 2024), which achieves state-of-the-art results on nanoGPT by orthogonalizing the update directions before applying them.
Pruning. Neural networks are often overparameterized, which can enhance generalization and avoid poor local minima. But such models then suffer from excessive memory and computational demands, making them less efficient for deployment in real-world applications (Chang et al., 2021). Pruning reduces the number of parameters, resulting in more compact and efficient models that are easier to deploy. A comprehensive review on pruning is beyond the scope of this paper due to space limitations and the diversity of approaches. We refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024) and the references therein for detailed reviews. Pruning by singular value thresholding has recently shown promising results, particularly in natural language processing (Chen et al., 2021), and is often used along with various enhancements such as importance weights and data whitening for effective compression of large language models (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024).
@@ 2 Matrix Factorization with a Diagonal Component
Consider the problem of reconstructing a matrix $\smash{X\in\mathbb{R}^{d\times c}}$ from a set of linear measurements $\smash{b=\mathcal{A}(X)\in\mathbb{R}^{n}}$. We focus on recovering a PSD matrix, without loss of generality since the general case can be reformulated as a PSD matrix sensing problem (Park et al., 2017). We particularly focus on the data-scarce setting where $n\ll d^{2}$, a notable example is matrix completion. This problem is inherently under-determined, but successful recovery is possible if $X$ is low-rank (Candes & Recht, 2012).
One popular approach is the BM factorization, which reparametrizes the decision variable $X$ as $UU^{\top}$ with $U\in\mathbb{R}^{d\times r}$, and $r$ is a positive integer that controls the rank of the resulting product. Then, the problem can be formulated as:
@@
While finding the global minimum is intractable, a local solution can be found by using gradient descent (Lee et al., 2016). Initializing at some $U_{0}\in\mathbb{R}^{d\times r}$, perform:
$U_{k+1}=U_{k}-\eta\nabla_{U}f(U_{k}U_{k}^{\top}),$
where $\eta>0$ is the step-size, and the gradient is computed as $\nabla_{U}f(UU^{\top})=2\nabla f(UU^{\top})U$.
Selecting the factorization rank $r$ is a critical decision. A small $r$ may lead to spurious local minima, resulting in inaccurate outcomes (Waldspurger & Waters, 2020). Conversely, a large $r$ might weaken rank regularization, rendering the problem underdetermined. Conventional wisdom suggests finding a moderate compromise between these two extremes. However, a key observation in (Gunasekar et al., 2017) is that the gradient method applied to (1) exhibits a tendency towards approximately low-rank solutions even when $r=d$.
@@ 2.1 The Proposed Factorization
$\min_{\begin{subarray}{c}U\in\mathbb{R}^{d\times r}\\ D\in\mathbb{R}^{r\times r}\end{subarray}}~\frac{1}{2}\|\mathcal{A}(UDU^{\top})-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\|U\|_{F}\leq\alpha,~~D_{ii}\geq 0,~~D_{ij}=0,\quad\forall i~\text{and}~\forall j\neq i$ (2)
@@
We perform simultaneous projected-gradient updates on $U$ and $D$ with step-size $\eta>0$:
@@
2.2 Experiments on Matrix Factorization
We present numerical experiments that compare the proposed approach with the classical BM factorization on synthetic matrix completion and on a real Fourier ptychography phase retrieval task.
2.2.1 Matrix Completion
We set up a synthetic matrix completion problem to recover a PSD matrix $\smash{X_{\natural}=U_{\natural}U_{\natural}^{\top}}$, where the entries of $\smash{U_{\natural}\in\mathbb{R}^{100\times 3}}$ are drawn independently from $N(0,1)$. We randomly sample $n=900$ entries of $X_{\natural}$ and store them in the vector $b\in\mathbb{R}^{n}$. The goal is to recover $X_{\natural}$ from $b$ by solving problems (1) and (2).
Although the ground-truth rank is three, we adopt a full-rank factorization with $r=d$ in order to study the implicit bias. For initialization, we generate $U_{0}\in\mathbb{R}^{d\times d}$ with entries drawn independently from $N(0,1)$; and rescale it to have Frobenius norm $\xi>0$, which sets its initial distance from the origin. For the UDU formulation, we set $D_{0}=I$ and the constraint parameter $\alpha=1$. Our experiments focus on how initialization ($\xi$) and step size ($\eta$) affect the singular value spectrum of the solutions.
We first investigate the impact of step-size. To this end, we fix $\xi=10^{-2}$ and test different values of $\eta$. The results are shown in Figure 1. In the left panel, we plot the objective residual as a function of iterations. As expected, we observe that a smaller step-size slows down convergence. The right panel presents the singular value spectra after $10^{6}$ iterations, revealing no clear correlation between step size and the spectral decay of the solutions. Notably, while the classical BM factorization finds an approximately low-rank solution, the proposed factorization converges to a solution of true rank three, matching the ground truth.
Impact of step-size ($\eta$), with fixed initialization.
Impact of initial distance to origin $(\xi)$, with fixed step-size.